OpenStudy (anonymous):

Solve for x: one over eight (8x + 15) = 24

jagr2713 (jagr2713):

is it \[\frac{ 1 }{ 8 }(8x+15)=24\]

OpenStudy (anonymous):

9 49 one hundred seventy-seven over eight two hundred seven over eight

OpenStudy (anonymous):

those are the given answers

jagr2713 (jagr2713):

What do you think the answer is or you dont know?

OpenStudy (anonymous):

i dont know what it is

jagr2713 (jagr2713):

Ok first If we simplify \[\frac{ 1 }{ 8 }(8x+15) \]what do we get?

OpenStudy (anonymous):

1/8 *23 ?

OpenStudy (anonymous):

1/8*23x

jagr2713 (jagr2713):

Nope we are just simply moving the 8 to the 8x+15 so we get\[\frac{ 8x+15 }{ 8 }=24\]

jagr2713 (jagr2713):

So what do we do next?

OpenStudy (anonymous):

add 8x + 15? sorry if i get this wrong i just dont understand it

jagr2713 (jagr2713):

Its ok to be wrong. Your learning now :D Ok no we do what with the 8 and 24

OpenStudy (anonymous):

add or divide it?

jagr2713 (jagr2713):

Nope whens its like that you multiply them other to bring it to a one line equation so we get?

OpenStudy (anonymous):

okay i get it now

jagr2713 (jagr2713):

So what do we get?

OpenStudy (anonymous):

8*24=192

jagr2713 (jagr2713):

where did u get the 24 from?

OpenStudy (anonymous):

ohh sorry i think i get 49?....

jagr2713 (jagr2713):

Ohh sorry yea 192 lol so we are left with 8x+15=192 Ok now can u solve from here

OpenStudy (anonymous):

subtract 15 from both sides, and divide both sides by 8?

OpenStudy (anonymous):

so it is C

jagr2713 (jagr2713):

Great job :D\[ \huge{\color{purple}{\textbf{W}} \color{orange}{\cal{E}} \color {green}{\mathbb{L}} \color{blue}{\mathsf{C}} \color{maroon}{\rm{O}} \color{red}{\tt{M}} \color{gold}{\tt{E}} \space \color{orchid} {\mathbf{T}} \color{Navy}{\mathsf{O}} \space \color{OrangeRed} {\boldsymbol{O}} \color{Olive}{\mathbf{P}} \color{Lime}{\textbf{E}} \color{DarkOrchid}{\mathsf{N}} \color{Tan}{\mathtt{S}} \color {magenta}{\mathbb{T}} \color{goldenrod}{\mathsf{U}} \color {ForestGreen}{\textbf{D}} \color{Salmon}{\mathsf{Y}} \ddot \smile }\] if you want to know more about this site, Here is a tutorial about this site : http://openstudy.com/study#/updates/5525c0fce4b030ceb3f357e8

OpenStudy (anonymous):

thank you!!!!

jagr2713 (jagr2713):

Anytime :D